• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jwexler

    Log entry time 16:01:57 on May17,2010

    Entry number 322975

    keyword=Shift summary - Day 5/17

    Shift summary:

    08:00 - MCC not ready to return beam yet, and EPICS not responding.
    08:15 - EPICS responding but the 'white' alarm is still on.
    08:30 - Recieving tune beam.
    08:50 - Recieved beam, checked raster, starting a production run.
    09:20 - Requested controlled access for dipole work.
    11:30 - Saw target vacuum is increasing.
    12:10 - Target vacuum at 2:10^-5. Yellow 'minor' alarm started.
    13:15 - Changed slug number to 24, IHWP OUT. Also, after changing the alarm settings, the yellow 'minor' vacuum alarm has triggered for vacuum at pressure 5*10^-5
    14:00 - Ed Folts added a second heater to try and preserve target vacuum. Should be removed before running production.
    15:50 - Jack fixed the vacuum. We're going to 10uA in order to try and restore vacuum via heating. We have left the second heater on the O-ring in order to try and assist that.
    15:55 - Beginning beam recovery. Going to empty target.

    RUN: 3916